# CMB Constraints on Natural Inflation with Gauge Field Production

### v1

@@ 1 Introduction
The minimal model with a periodic potential for a pNGB [11, 12] of natural inflation is in tension with the recent data [1], especially for the sub-Planckian values of the symmetry-breaking scale $f\lesssim M_{{}_{\mathrm{Pl}}}$ [^2] Throughout the draft, we will be using the reduced Planck mass as $M_{{}_{\mathrm{Pl}}}$, and if not written explicitly, it will be set to $M_{{}_{\mathrm{Pl}}}=1$.. The natural inflation model with a canonical kinetic energy term is ruled out (at $2$-$\sigma$) for all values of $f$ when the Planck data is combined with recent BICEP/Keck and BAO data; see Fig. 5 of reference [13]. The question that naturally arises is whether the model can be rescued by keeping its basic features intact. In this article, we would like to address this issue. The symmetry-breaking scale $f$ being larger than the Planck mass means that the global symmetry is broken above the quantum gravity scale where the validity of effective field theory is on shaky ground [14, 15]. At the same time, it has been conjectured that it is difficult to realize super-Planckian axion decay constant in any fundamental theory like in String Theory [16]. However, there exist models of many fields with sub-Planckian decay constants for an individual field, but effectively the models behave as one with an effective super-Planckian decay constant [17, 18, 9, 10, 19, 20]. Note that the CMB observables are sensitive only to this effective decay constant. The revised question that arises is that even for super-Planckian $f_{\rm eff}$, when is the model consistent with the data?
[…]
@@
Our goal is to see whether the gauge field production in natural inflation with periodic potential can keep the model alive. The effects of gauge field can have broadly three regimes. Obviously, in one case, the effects can be so small (small $\alpha$) that it does not affect background and perturbations both at the CMB scales and small scales. In the second case, the production of gauge fields is very high to start with and therefore affects both background and perturbations throughout the evolution of the inflaton [48, 49, 23]. Finally, the effects can be such that the gauge field affects the perturbations at the CMB scales, but not the background. But, due to the exponential growth of gauge fields, it affects the evolution at the later stages of inflation. We broadly work in this limit.  Therefore, in our case, the effects of back-reaction are small at $60$ e-foldings before the end of inflation, but we reach the strong back-reaction regime at a later stage. We will conclude that the model can be consistent only with the super-Planckian effective decay constant, and in that case also, only for a range of the coupling constant $\alpha$.
@@ 2 Pseudoscalar inflation with a gauge field
In this section, we discuss pseudoscalar inflaton $\phi$ coupled to an Abelian gauge field [22]. In particular, we will focus on the case of natural inflation where the inflaton potential is periodic [11, 12]. In this case, a higher-dimensional coupling between the inflaton and the gauge field introduces an instability in the theory, and it leads to an exponential production of the gauge fields [21, 22]. It is well known that this coupling induces a back-reaction both on the background dynamics [23, 24] and in the the perturbations [38, 27, 25, 24, 26], leading to several observational consequences.
@@ 2.1 Gauge field production
where $\xi=\frac{\alpha\,\dot{\phi_{0}}}{2\,f\,H}$. From the above equation, we see that one of the two modes grows exponentially if that mode satisfies the condition $k/a\,H<2|\xi|$. In another way, Eq. (2.12) describes a tachyonic instability for the $A_{k}^{+}$ mode when we consider the initial velocity of the inflaton field to be positive. In this case, $\xi$ is always positive and increases on average till the end of inflation. During the initial stages of inflation, the value of $\xi$ remains nearly constant due to the slow-roll condition. Assuming $\xi$ to be constant, we can write down the solution of Eq. (2.12) for the modes, $(8\,\xi)^{-1}\lesssim k/(aH)\lesssim 2\,\xi$. Additionally, we can also write write down $\langle\mathbf{E}.\mathbf{B}\rangle$, and $\frac{1}{2}\langle\mathbf{E}^{2}+\mathbf{B}^{2}\rangle$ in terms of $A_{k}^{\pm}$  [23, 25, 26]:
@@ 2.2 Back-reaction of gauge fields on inflation dynamics
In Fig. 2, the evolution of the first slow-roll parameter $\epsilon_{H}=-\dot{H}/H^{2}$ and the Hubble parameter are shown for various values of the coupling parameter $\alpha$. Inflation ends when $\epsilon_{H}=1$ and these are marked by the vertical lines. Several stars correspond to the epoch of $60$ $e$-foldings before the end of inflation. From the $\epsilon_{H}$ plot, we observe that the presence of the gauge field prolongs inflation duration since the gauge field decelerate the dynamics of the inflaton field. In the right panel of Fig. 2, we have also depicted the evolution of the Hubble parameter with time. In Fig. 3, the inflaton dynamics and the phase diagram have been shown. From these figures, we can conclude that the initial inflaton dynamics mimic the usual cold inflationary behavior.
@@ 3 Scalar and tensor power spectra
It should be mentioned here that the above expression accounts for the contributions coming from the source term only. The total power spectra including the quantum vacuum fluctuations is
@@ 4 Imprints on CMB observables
In practice, we evaluate the scalar power spectrum as a function of the number of e-foldings $N$ using the values of background quantities appearing in the Eq. (3.8). Note that the effects of back-reaction are incorporated in the background quantities. We subsequently convert the spectrum in terms of comoving momenta $k$ to produce the left panel of Fig. 4 where scalar spectrum $P_{\zeta}$ is plotted against $k$. In the plot, we show the scalar power spectrum for different values of $\alpha$ for the fixed value of $f=6$. The red line corresponds to the usual slow-roll inflation in the absence of a gauge field. The scalar power spectrum shows several interesting features for $\alpha\neq 0$. The most prominent one is the sharp rise in the value of the spectrum at small scales. The rise of the spectrum happens for smaller $k$ values (larger wavelength) for larger $\alpha$. Also, due to the strong back-reaction at the later stage of inflation, the power spectrum shows oscillatory features at small scales. It is essential to reemphasize that the spectrum does not diverge at smaller scales due to the consideration of the back-reaction effect in calculations of both the scalar power spectrum and $\xi$. In this section, we focus on the power spectrum at the CMB scales. In the plot, we also show several existing and future observational constraints on the scalar power spectrum from Planck [1], Lyman-alpha [59], PIXIE [60], COBE/Firas [61], and PTA [62].
@@
Figure 5: The scalar spectral index $n_{s}$ vs the tensor-to-scalar ratio $r$ for various values of the coupling constant $\alpha$. The continuous purple line corresponds to the case for $\alpha=0$ as we vary the value of $f$. The orange, red, and green dots of different colors correspond to different values of $f$ for $N_{*}=60$. On the other hand, black diamond corresponds to $N_{*}=50$ and $f=5$. The bigger contour shows the $2$-$\sigma$ constraints coming from only Planck data, and the smaller contour shows the $2$-$\sigma$ constraints when the Planck data [1] is combined with BK18 and BAO [13].
@@
and taking the ratio $P_{\zeta}(k)/P_{h}(k)$ at the pivot scale. In Fig. 5, we show the model predictions for $n_{s}$ and $r$ as we vary $\alpha$ for several values of $f$. We also show $2$-$\sigma$ allowed contours for two different observational data sets: 2018 Planck TT, TE, EE+lowE+lensing [1] (henceforth will be mentioned as Planck18) and 20018 Planck TT, TE, EE+lowE+lensing+BK18+BAO [13] (henceforth will be called Planck18+BK18). The solid purple line in Fig. 5 shows the model predictions $(N_{*}=60)$ for canonical pNGB potential when $f$ is continuously varied. First of all, the model is consistent at $2$-$\sigma$ for $f\gtrsim 5.5$ only when the Planck18 data is considered. On the other hand, the model is excluded at $2$-$\sigma$ for all values of $f$ when the Planck18 data is combined with the BK18 and BAO data. In this sense, the pNGB inflation model with periodic potential is under tension with the latest observational data. Note that reducing the value of $N_{*}$ makes the situation worse.
To understand the effects of the gauge field production we will first discuss the case of $f=5$ with $N_{*}=60$ case as we vary $\alpha$. These variations are depicted by red dot points in Fig. 5. As we increase the value of the coupling constant $\alpha$ from zero, the spectral index $n_{s}$ starts to decrease, and predictions go further away from the $2$-$\sigma$ contour given by the Planck18 data. But, interestingly, the value of the spectral index makes a turn and then it starts to increase. At some point, it becomes consistent with the latest Planck18 data, but it still remains outside the contour for Planck, BK18, and BAO combined data. In this case, the spectral index is consistent for $\alpha$ between 300 and 313 (see the red dots in Fig.5) when only Planck18 data is considered. Therefore, a coupling between the gauge field and the pseudo-scalar inflaton helps to alleviate the tension of the pNGB model and makes it consistent in the $n_{s}$-$r$ plane.
@@
The turn of the spectral index happens due to a competition between the vacuum contribution and the gauge field contribution to the scalar spectrum. For the vacuum part of the scalar power spectrum, the slope decreases at the pivot scale as the coupling constant $\alpha$ increases. This decrease in slope occurs because the pivot scale shifts, which is essentially caused by an increase in the duration of inflation. However, the slope of the scalar power spectrum due to the contributions of the gauge field increases with the coupling constant $\alpha$. At larger values of $\alpha$ the spectral index even becomes greater than $1$ where the gauge fields mainly source the scalar power spectrum [23]. This can be well understood in Fig. 6 where we show the amplitude of the power spectrum only around the pivot scale. We see that the slope makes a turn as we increase $\alpha$. To make this plot, we have adjusted the amplitude at the pivot scale to be the same for all values of $\alpha$ by changing the value $\Lambda$. For the allowed range of $300\lesssim\alpha\lesssim 313$, $r$ does not change appreciably and remains within the observational limit.
As mentioned before, the gauge field not only affect the $n_{s}$ and $r$, it also induce non-gaussianities and running in the spectrum. Therefore, it is important to check whether the above-mentioned allowed range for $\alpha$ is consistent with the observational limits on running of the spectral index and the non-gaussianities. In the left panel of Fig. 7, we show the running of the spectral index for different values of the coupling constant $\alpha$ for $f=4,~{}5$. We observe that the running of the spectral index grows with $\alpha$ due to the growing contributions from the gauge field. We also show the $2\sigma$ limit on the running $dn_{s}/d\ln k=0.002\pm 0.020$ coming from the Planck18 data [1]. In particular, we find that the allowed range of $\alpha$ in the $n_{s}$-$r$ plane is also within the observational bound provided by the running of the spectral index. Therefore, we find that the values of $\alpha$ for which the model is consistent with $n_{s}$-$r$ observations are also consistent with running. Note that we have plotted the running of the spectrum up to certain values of $\alpha$ as larger values of $\alpha$ are excluded only by the $n_{s}$ constraints.
Figure 7: The left panel shows the running of scalar spectral index vs the coupling constant $\alpha$ for $f=4,~{}5$. The shaded regions are excluded by Planck18 data [1]. The right panel illustrates the non-causality parameter vs $\alpha$ for $f=4,~{}5$. The horizontal line shows the upper limit from observations [63].
@@
for $\xi\gg 1$. Now, we plug the numerically evaluated values of $P_{\zeta}(k)_{\rm vac}$, $P_{\zeta}(k)$ and $\xi$ at the pivot scale to get the value of $f^{\rm equil}_{\rm NL}$. It is shown in the right panel of Fig. 7 for two values of $f$. The orange dotted line shows the predicted non-gaussianities for different values of $\alpha$ for $f=5$. The horizontal line is the observational upper limit $f_{\rm NL}^{\rm equil}=-26\pm 94$ [63]. We see that all relevant values of $\alpha$ that are consistent with $n_{s}$ are also consistent with the non-gaussianity bound. We can summarise the result for $f=5$ by noting that the values $\alpha$ allowed by $n_{s}$-$r$ constraint, i.e. $300\lesssim\alpha\lesssim 313$, are also consistent with running and non-gausianities. The constraint on $\alpha$ can be translated to the value $\xi$ at the CMB scale, and for $f=5$, it is $2.23\lesssim\xi_{*}\lesssim 2.36$.
The constraints on $\alpha$ and therefore $\xi_{*}$ are summarised in Table. 1 for several values of $f$ and $N_{*}$ when only Planck18 data is considered. We see that for $f=6$, there is only an upper limit on $\alpha=305$. In fact, the existence of an upper limit continues for larger values of $f$ also as long as the corresponding $r$ does not become too large. On the other hand for $f\lesssim 5.5$, the model is consistent only for a range of $\alpha$ or $\xi_{*}$. If $f$ becomes smaller than the Planck mass, it is difficult to achieve successful inflation of the required duration of inflation. When we consider $N_{*}=50$ (f = 5), the required value of $\alpha$ to make the model consistent with the data is smaller. In the case when the Planck data is combined with BK18 and BAO data the model is consistent only for $f=4$ and $347\lesssim\alpha\lesssim 352$ with corresponding values of $2.38\lesssim\xi_{*}\lesssim 2.44$. This happens mainly due to the improved limit on the tensor-to-scalar ratio when two data sets are combined.
@@
Table 1: Constraints on $\alpha$ and $\xi_{*}$ for several values of $f$ and $N_{*}$ when only Planck data is used.

### v2

@@ 1 Introduction
The minimal model with a periodic potential for a pNGB [11, 12] of natural inflation is in tension with the recent data [1], especially for the sub-Planckian values of the symmetry-breaking scale $f\lesssim M_{{}_{\mathrm{Pl}}}$ [^2] Throughout the draft, we will be using the reduced Planck mass as $M_{{}_{\mathrm{Pl}}}$, and if not written explicitly, it will be set to $M_{{}_{\mathrm{Pl}}}=1$.. The natural inflation model with a canonical kinetic energy term is ruled out (at $2$-$\sigma$) for all values of $f$ when the Planck data is combined with recent BICEP/Keck and BAO data; see Fig. 5 of reference [13]. The question that naturally arises is whether the model can be rescued by keeping its basic features intact. In this article, we would like to address this issue. The symmetry-breaking scale $f$ being larger than the Planck mass means that the global symmetry is broken above the quantum gravity scale where the validity of effective field theory is on shaky ground [14, 15, 16]. At the same time, it has been conjectured that it is difficult to realize super-Planckian axion decay constant in any fundamental theory like in String Theory [17]. However, there exist models of many fields with sub-Planckian decay constants for an individual field, but effectively the models behave as one with an effective super-Planckian decay constant [18, 19, 9, 10, 20, 21]. Note that the CMB observables are sensitive only to this effective decay constant. The revised question that arises is that even for super-Planckian $f_{\rm eff}$, when is the model consistent with the data?
Other than the self-coupling, it is easily conceivable that the inflaton field is also coupled with the matter fields present in the theory. The pNGB field, due to the parity-violating nature (a pseudo-scalar), is generally coupled with the gauge field via the 5-dimensional operator $\mathcal{L}_{\rm int}\supset-\frac{\alpha}{4\,f}\phi\,F\,\tilde{F}$, where $\alpha$ is the dimensionless coupling strength, $F(\tilde{F})$ is the gauge field (dual) tensors associated with some $U(1)$ gauge field $A_{\mu}$ and $f$ is the symmetry breaking scale. The coupling term allows inflaton energy to dissipate into the gauge fields and one of the helicity modes of the gauge field gets excited during inflation [22, 23, 24, 25]. This excited mode grows exponentially depending on the inflaton velocity dependent parameter $\xi(t)\equiv\frac{\alpha\,\dot{\phi_{0}}}{2fH}$, where $\phi_{0}$ is the homogenous part of the inflaton field. As the velocity of the inflaton increases, the gauge fields are produced copiously and that starts to back react to the inflation dynamics leading to a prolonged duration of inflation [26, 27]. The production of gauge field also starts to affect the scalar perturbations via inverse decay of the gauge field, and the spectrum becomes highly non-gaussian at the CMB scales for $\xi_{*}\gtrsim 1$ [28, 29]. Throughout the paper, the quantities with the subscript * indicate values at the CMB scale.
@@
Our goal is to see whether the gauge field production in natural inflation with periodic potential can keep the model alive. The effects of gauge field can have broadly three regimes. Obviously, in one case, the effects can be so small (small $\alpha$) that it does not affect background and perturbations both at the CMB scales and small scales. In the second case, the production of gauge fields is very high to start with and therefore affects both background and perturbations throughout the evolution of the inflaton [51, 52, 26, 62]. Finally, the effects can be such that the gauge field affects the perturbations at the CMB scales, but not the background. But, due to the exponential growth of gauge fields, it affects the evolution at the later stages of inflation. We broadly work in this limit. Therefore, in our case, the effects of back-reaction are small at $60$ e-foldings before the end of inflation, but we reach the strong back-reaction regime at a later stage. We will conclude that the model can be consistent only with the super-Planckian effective decay constant, and in that case also, only for a range of the coupling constant $\alpha$.
@@ 2 Pseudoscalar inflation with a gauge field
In this section, we discuss pseudoscalar inflaton $\phi$ coupled to an Abelian gauge field [23]. In particular, we will focus on the case of natural inflation where the inflaton potential is periodic [11, 12]. In this case, a higher-dimensional coupling between the inflaton and the gauge field introduces an instability in the theory, and it leads to an exponential production of the gauge fields [22, 23]. It is well known that this coupling induces a back-reaction both on the background dynamics [26, 27] and in the perturbations [41, 30, 28, 27, 29], leading to several observational consequences.
@@ 2.1 Gauge field production
where $\xi=\frac{\alpha\,\dot{\phi_{0}}}{2\,f\,H}$. From the above equation, we see that one of the two modes grows exponentially if that mode satisfies the condition $k/a\,H<2|\xi|$. In another way, Eq. (2.12) describes a tachyonic instability for the $A_{k}^{+}$ mode when we consider the initial velocity of the inflaton field to be positive. In this case, $\xi$ is always positive and increases on average till the end of inflation. During the initial stages of inflation, the value of $\xi$ remains nearly constant due to the slow-roll condition. Assuming $\xi$ to be constant, we can write down the solution of Eq. (2.12) for the modes, $(8\,\xi)^{-1}\lesssim k/(aH)\lesssim 2\,\xi$. Additionally, we can also write down $\langle\mathbf{E}.\mathbf{B}\rangle$, and $\frac{1}{2}\langle\mathbf{E}^{2}+\mathbf{B}^{2}\rangle$ in terms of $A_{k}^{\pm}$  [26, 28, 29]:
@@ 2.2 Back-reaction of gauge fields on inflation dynamics
In Fig. 2, the evolution of the first slow-roll parameter $\epsilon_{H}=-\dot{H}/H^{2}$ and the Hubble parameter are shown for various values of the coupling parameter $\alpha$. Inflation ends when $\epsilon_{H}=1$ and these are marked by the vertical lines. Several stars correspond to the epoch of $60$ $e$-foldings before the end of inflation. From the $\epsilon_{H}$ plot, we observe that the presence of the gauge field prolongs inflation duration since the gauge field decelerate the dynamics of the inflaton field. In the right panel of Fig. 2, we have also depicted the evolution of the Hubble parameter with time. In Fig. 3, the inflaton dynamics and the phase diagram have been shown. From these figures, we can conclude that the initial inflaton dynamics mimic the usual cold inflationary behavior. 444 The inclusion of gradient terms or spatial inhomogeneities in the inflaton field could affect the background dynamics, such as the duration of inflation, the Hubble parameter, and the slow-roll parameters, as shown in Fig. 2 and Fig. 3. However, we believe that at the background level, these changes are not substantial enough to alter the overall scenario significantly (see [59]).
@@ 3 Scalar and tensor power spectra
It should be mentioned here that the above expression accounts for the contributions coming from the source term only. It should be noted that while evaluating the expression (3.7), spatial inhomogeneities were not taken into account. Including such terms could modify the behavior of the power spectrum at small scales. However, these small-scale modifications do not significantly impact the CMB scales. The total power spectra including the quantum vacuum fluctuations is
@@ 4 Imprints on CMB observables
In practice, we evaluate the scalar power spectrum as a function of the number of e-foldings $N$ using the values of background quantities appearing in the Eq. (3.8). Note that the effects of back-reaction are incorporated in the background quantities. However, as previously mentioned, the spatial inhomogeneities of the inflaton field have not been taken into account. We subsequently convert the spectrum in terms of comoving momenta $k$ to produce the left panel of Fig. 4 where scalar spectrum $P_{\zeta}$ is plotted against $k$. In the plot, we show the scalar power spectrum for different values of $\alpha$ for the fixed value of $f=6$. The red line corresponds to the usual slow-roll inflation in the absence of a gauge field. The scalar power spectrum shows several interesting features for $\alpha\neq 0$. The most prominent one is the sharp rise in the value of the spectrum at small scales. The rise of the spectrum happens for smaller $k$ values (larger wavelength) for larger $\alpha$. Also, due to the strong back-reaction at the later stage of inflation, the power spectrum shows oscillatory features at small scales. It is essential to reemphasize that the spectrum does not diverge at smaller scales due to the consideration of the back-reaction effect in calculations of both the scalar power spectrum and $\xi$. In this section, we focus on the power spectrum at the CMB scales. In the plot, we also show several existing and future observational constraints on the scalar power spectrum from Planck [1], Lyman-alpha [63], PIXIE [64], COBE/Firas [65], and PTA [66].
@@
Figure 5: The scalar spectral index $n_{s}$ vs the tensor-to-scalar ratio $r$ for various values of the coupling constant $\alpha$. The continuous purple line corresponds to the case for $\alpha=0$ as we vary the value of $f$. The orange, red, and green dots of different colors correspond to different values of $f$ for $N_{*}=60$. On the other hand, black diamond corresponds to $N_{*}=50$ and $f=5$. The grey colored contour shows the $2$-$\sigma$ constraints coming from only Planck18 data, and the blue contours show the $1$-$\sigma$ (darker) and $2$-$\sigma$ (lighter) constraints when the Planck18 data [1] is combined with BK18 and BAO [13].
@@
and taking the ratio $P_{\zeta}(k)/P_{h}(k)$ at the pivot scale. In Fig. 5, we show the model predictions for $n_{s}$ and $r$ as we vary $\alpha$ for several values of $f$. We also show $2$-$\sigma$ allowed contours for 2018 Planck TT, TE, EE+lowE+lensing [1] (henceforth will be mentioned as Planck18) and $1$-$\sigma$ and $2$-$\sigma$ contours for 2018 Planck TT, TE, EE+lowE+lensing+BK18+BAO [13] (henceforth will be called Planck18+BK18). The solid purple line in Fig. 5 shows the model predictions $(N_{*}=60)$ for canonical pNGB potential when $f$ is continuously varied. First of all, the model is consistent at $2$-$\sigma$ for $f\gtrsim 5.5$ only when the Planck18 data is considered. On the other hand, the model is excluded at $2$-$\sigma$ for all values of $f$ for Planck18+BK18 data. In this sense, the pNGB inflation model with periodic potential is under tension with the latest observational data. Note that reducing the value of $N_{*}$ makes the situation worse.
To understand the effects of gauge field production, we first focus on the interesting case with $f=4$ and $N_{*}=60$, as we vary $\alpha$. These variations are shown by the orange dotted points in Fig. 5 where $\alpha=0$ is well outside the Planck18 $2$-$\sigma$ contour. It is noted above, for $\alpha=0$, the model is excluded by Planck18 data at $2$-$\sigma$. As we increase the value of the coupling constant $\alpha$ from zero, the spectral index $n_{s}$ initially decreases, moving further away from the 2-$\sigma$ contour provided by the Planck18 data. Interestingly, the value of the spectral index makes a turn and then it starts to increase as the value of $\alpha$ is raised. At some point, it becomes consistent with the latest Planck18 data. If we continue increasing $\alpha$, the predictions eventually enter $2$-$\sigma$ followed by $1$-$\sigma$ contours even for the combined data set Planck18 + BK18. In this case, the spectral index remains consistent with the combined data set for $347\lesssim\alpha\lesssim 352$ at the $2$-$\sigma$ level and for $349\lesssim\alpha\lesssim 351$ at the $1$-$\sigma$ level (see the orange dots in Fig.5). Therefore, a coupling between the gauge field and the pseudo-scalar inflaton helps to alleviate the tension of the pNGB model and makes it consistent in the $n_{s}$-$r$ plane. In summary, natural inflation with $\alpha=0$ is very strongly disfavored by Planck18+BK18 data, irrespective of the value of $f$. As an example, we see that for $f=4$, a narrow range of $\alpha$ brings the model to agree with Planck+Keck.
The above-mentioned trend of $n_{s}$ dependence on $\alpha$ remains valid for higher values of $f$- see red and green dots of $f=5$ and $f=6$ respectively. In the case of $f=5$, the model becomes consistent ($2$-$\sigma$) with only Planck18 data for some range of $\alpha$ values, but it never becomes consistent with Planck18+BK18 data at $2$-$\sigma$. For the case of $f=6$, the model remains consistent ($2$-$\sigma$) with Planck18 data for $\alpha\lesssim 305$.
@@
The turn of the spectral index happens due to a competition between the vacuum contribution and the gauge field contribution to the scalar spectrum. For the vacuum part of the scalar power spectrum, the slope decreases at the pivot scale as the coupling constant $\alpha$ increases. This decrease in slope occurs because the pivot scale shifts, which is essentially caused by an increase in the duration of inflation. However, the slope of the scalar power spectrum due to the contributions of the gauge field increases with the coupling constant $\alpha$. At larger values of $\alpha$ the spectral index even becomes greater than $1$ where the gauge fields mainly source the scalar power spectrum [26]. This can be well understood in Fig. 6 where we show the amplitude of the power spectrum only around the pivot scale. We see that the slope makes a turn as we increase $\alpha$. To make this plot, we have adjusted the amplitude at the pivot scale to be the same for all values of $\alpha$ by changing the value $\Lambda$. For the allowed range of $\alpha$, $r$ does not change appreciably and remains within the observational limit for the combined data.
As mentioned before, the gauge field not only affect the $n_{s}$ and $r$, it also induce non-gaussianities and running in the spectrum. Therefore, it is important to check whether the above-mentioned allowed range for $\alpha$ is consistent with the observational limits on running of the spectral index and the non-gaussianities. In the left panel of Fig. 7, we show the running of the spectral index for different values of the coupling constant $\alpha$ for $f=4,~{}5$ and $6$. We observe that the running of the spectral index grows with $\alpha$ due to the growing contributions from the gauge field. We also show the $2\sigma$ limit on the running $dn_{s}/d\ln k=0.002\pm 0.020$ coming from the Planck18 data [1]. In particular, we find that the allowed range of $\alpha$ in the $n_{s}$-$r$ plane is also within the observational bound provided by the running of the spectral index. Therefore, we find that the values of $\alpha$ for which the model is consistent with $n_{s}$-$r$ observations are also consistent with running. Note that we have plotted the running of the spectrum up to certain values of $\alpha$ as larger values of $\alpha$ are excluded only by the $n_{s}$ constraints.
Figure 7: The left panel shows the running of scalar spectral index vs the coupling constant $\alpha$ for $f=4,~{}5~{}\text{and}~{}6$. The shaded regions are excluded by Planck18 data [1]. The right panel illustrates the non-causality parameter vs $\alpha$ for $f=4,~{}5~{}\text{and}~{}6$. The horizontal line shows the upper limit from observations [67].
@@
for $\xi\gg 1$. Now, we plug the numerically evaluated values of $P_{\zeta}(k)_{\rm vac}$, $P_{\zeta}(k)$ and $\xi$ at the pivot scale to get the value of $f^{\rm equil}_{\rm NL}$. It is shown in the right panel of Fig. 7 for three values of $f$. The orange dotted line shows the predicted non-gaussianities for different values of $\alpha$ for $f=4$. The horizontal line is the observational upper limit $f_{\rm NL}^{\rm equil}=-26\pm 94$ [67]. We see that all relevant values of $\alpha$ that are consistent with $n_{s}$ are also consistent with the non-gaussianity bound. We can summarise the result for $f=4$ by noting that the values $\alpha$ allowed by $n_{s}$-$r$ constraint, i.e. $347\lesssim\alpha\lesssim 352$ ($2$-$\sigma$ Planck18+BK18), are also consistent with running and non-gausianities. The constraint on $\alpha$ can be translated to the value $\xi$ at the CMB scale, and for $f=4$, it is $2.38\lesssim\xi_{*}\lesssim 2.44$.
The constraints on $\alpha$ and therefore $\xi_{*}$ are summarised in Table. 1 for several values of $f$ and $N_{*}$ when only Planck18 data is considered. For the case where $f=5$ and $N_{*}=60$, we find that the model is consistent only with Planck18 data at the 2-$\sigma$ level for the parameter values in the range $300\lesssim\alpha\lesssim 313$. The corresponding values for $\xi_{*}$ fall within the range $2.32\lesssim\xi_{*}\lesssim 2.44$. We see that for $f=6$, there is only an upper limit on $\alpha=305$. In fact, the existence of an upper limit continues for larger values of $f$ also as long as the corresponding $r$ does not become too large. On the other hand for $f\lesssim 5.5$, the model is consistent only for a range of $\alpha$ or $\xi_{*}$. If $f$ becomes smaller than the Planck mass, it is difficult to achieve successful inflation of the required duration of inflation. When we consider $N_{*}=50$ (f = 5), the required value of $\alpha$ to make the model consistent with the data is smaller. This happens mainly due to the improved limit on the tensor-to-scalar ratio when two data sets are combined.
@@
Table 1: Constraints on $\alpha$ and $\xi_{*}$ for several values of $f$ and $N_{*}$ when only Planck18 data is used.